on a standardized test, correct answers are worth 2 points and incorrect answers are worth -1/4 point. alexandra wants to score at least 90 points on the test. if there are 50 questions on the test, what is the greatest number of questions she can miss
Answer by richwmiller(17219)

x+y=50
2x-y/4=90
y=40/9
y=4 +
so y=4
max that can be missed is 4
If she misses 4 she loses 1 point and doesn't get 8
So she gets 91
check
46*2=92
92-1=91